# A Fourier-Jacobi Dirichlet series attached to modular forms of $\textup{SO}(2,4)$

### v1

@@ 8. Representation Numbers of Binary Hermitian Forms
The aim of this Section is to relate the expressions $n(\xi_{i};d)$ of equation (13) with $\xi_{i}\in L_{0}^{+}[\ell,2^{-1}\mathbb{Z}]/\widehat{\Gamma}(L_{0})$ to representation numbers of binary Hermitian quadratic forms. We recall,
@@
$n(\xi_{i};d)=\#\left\{s\in\mathbb{Z}^{2}/dS\mathbb{Z}^{2}\mid D\equiv\frac{1}{2}qS^{-1}[s]\pmod{qd},\,\,\left(\frac{\frac{1}{2}qS^{-1}[s]-D}{qd},S^{-1}s,d\right)^{t}=\gamma\xi_{i},\right.\\ \left.\gamma\in\widehat{\Gamma}^{+}(L_{0})\right\},$
where $D=-\ell q$, with $\ell$ a positive integer and $q$ the level of $S$. From Lemma 5.2 and Proposition 5.3, we have seen the identification
@@
induced by the map $f_{\omega}$. We set $f_{i}:=f_{\omega}^{-1}(\xi_{i})\in\Lambda^{+}[\ell,2^{-1}\mathbb{Z}]$, for $i=1,\ldots,h_{1}$. We first prove the following Lemma.
@@
Let $\Lambda=M_{2}(\mathcal{O}_{K})\cap U$, as above. We have the following two statements:
@@
The first statement follows from [Shi04, Lemma 6.2, (3)], and the second statement is [Shi06b, Lemma 4.5]. ∎
By this Lemma, we have in particular that $f_{i}\in\Lambda$ whenever $\ell\in\mathbb{Z}$. For a $d\in\mathbb{N}$, $f\in\Lambda$ with $\textup{det}(f)=\ell$, we set
$r(f;d)=\#\left\{h\in\mathcal{O}_{K}/d\mathcal{O}_{K}\mid h\overline{h}\equiv-\ell\pmod{d}\textup{ and }\begin{pmatrix}(h\overline{h}+\ell)/d&h\\ \overline{h}&d\end{pmatrix}\sim f\right\},$ (31)
where the notation $X\sim Y$ for two $2\times 2$ matrices $X$ and $Y$ means that there is a matrix $A\in\textup{SL}_{2}(\mathcal{O}_{K})$ such that $AX\overline{A}^{t}=Y$. We now prove the following Proposition.
@@
Let $\ell\in\mathbb{Z}_{>0}$ and $\xi_{i}\in L_{0}^{+}[\ell,2^{-1}\mathbb{Z}]$. Recall $f_{i}:=f_{\omega}^{-1}(\xi_{i})\in\Lambda^{+}[\ell,2^{-1}\mathbb{Z}]$. Then, for any $d\in\mathbb{N}$, we have $n(\xi_{i};d)=r(f_{i};d)$.
@@
We consider two cases:
Case 1: $m\equiv 1,2\pmod{4}$. In this case, from (16), $S=\textup{diag}(2,2m)$, and so we have $q=4m$ and $D=-4\ell m$. Therefore, if $s=(s_{1},s_{2})$, the congruence $D\equiv\frac{1}{2}qS^{-1}[s]\pmod{qd}$ can be written as
$4m\ell\equiv ms_{1}^{2}+s_{2}^{2}\pmod{4md}$,
with $s_{1}\pmod{2d},\,\,s_{2}\pmod{2md}$. Now, $m\mid s_{2}^{2}\implies m\mid s_{2}$ because $m$ is square-free. If we write $s_{2}\longmapsto ms_{2}$, the congruence becomes
$s_{1}^{2}+ms_{2}^{2}\equiv-4\ell\pmod{4d}$,
with $s_{1},s_{2}\pmod{2d}$. But since $m\equiv 1,2\pmod{4}$, $-m$ is not a quadratic residue $\pmod{4}$, so $s_{1}s_{2}$ must be both even. Hence, the congruence is equivalent to
$s_{1}^{2}+ms_{2}^{2}\equiv-\ell\pmod{d}$,
with $s_{1},\ s_{2}\pmod{d}$. Hence, by taking $h:=s_{1}+\omega s_{2}$, we see that this congruence is equivalent to $h\overline{h}\equiv-\ell\pmod{d}$ and $h\in\mathcal{O}_{K}/d\mathcal{O}_{K}$. Therefore, the map
$(s_{1},s_{2})\longmapsto h:=\frac{s_{1}}{2}+\omega\frac{s_{2}}{2m},$
induces a bijection of
$\left\{s\in\mathbb{Z}^{2}/dS\mathbb{Z}^{2}\mid D\equiv\frac{1}{2}qS^{-1}[s]\pmod{qd},\,\,\left(\frac{\frac{1}{2}qS^{-1}[s]-D}{qd},S^{-1}s,d\right)^{t}=\gamma\xi_{i},\right.\\ \left.\gamma\in\widehat{\Gamma}^{+}(L_{0})\right\}$
and the set
$\left\{h\in\mathcal{O}_{K}/d\mathcal{O}_{K}\mid h\overline{h}\equiv-\ell\pmod{d}\textup{ and }\begin{pmatrix}(h\overline{h}+\ell)/d&h\\ \overline{h}&d\end{pmatrix}\sim f_{i}\right\}.$
Hence $n(\xi_{i};d)=r(f_{i};d)$.
Case 2: $m\equiv 3\pmod{4}$. In this case, $q=m$ and hence $D=-\ell m$. We use the method used in [Psy25, Section 9.3]. In particular, if
$P=\begin{pmatrix}0&1\\ -1&2\end{pmatrix},\,\,\,Q=\begin{pmatrix}1&-(m+1)/2\\ 0&1\end{pmatrix}$,
we have $PSQ=\textup{diag}(1,m)$. Now, we set $t=Ps$ and the congruence $D\equiv\frac{1}{2}qS^{-1}[s]\pmod{qd}$ is equivalent to
$\frac{1}{2}qS^{-1}[P^{-1}t]\equiv D\pmod{qd}$,
where $t=(t_{1},t_{2})$ with $t_{1}\pmod{d}$ and $t_{2}\pmod{md}$. By substituting and performing the operations, we get
$mt_{1}^{2}-mt_{1}t_{2}+\frac{m+1}{4}t_{2}^{2}\equiv-\ell m\pmod{md}$,
with $t_{1},t_{2}$ as before. But $\gcd(m,(m+1)/4)=1$, so $m\mid t_{2}^{2}\implies m\mid t_{2}$, because $m$ is square-free. Hence, the congruence is equivalent to
$t_{1}^{2}-mt_{1}t_{2}+\frac{m(m+1)}{4}t_{2}\equiv-\ell\pmod{d}$,
with $t_{1},t_{2}\pmod{d}$. Hence, by taking $\displaystyle{h:=\left(t_{1}-\frac{m+1}{2}t_{2}\right)+\omega t_{2}}$ (recall that $\omega=(1+\sqrt{-m})/2$ in this case), we see that this is equivalent to $h\overline{h}\equiv-\ell\pmod{d}$ and $h\in\mathcal{O}_{K}/d\mathcal{O}_{K}$. Therefore, by changing back the variables, the map
${(s_{1},s_{2})\longmapsto h:=\left(\frac{m+1}{2}\frac{s_{1}}{m}-\frac{s_{2}}{m}\right)+\omega\frac{-s_{1}+2s_{2}}{m}}$
induces a bijection of
$\left\{s\in\mathbb{Z}^{2}/dS\mathbb{Z}^{2}\mid D\equiv\frac{1}{2}qS^{-1}[s]\pmod{qd},\,\,\left(\frac{\frac{1}{2}qS^{-1}[s]-D}{qd},S^{-1}s,d\right)^{t}=\gamma\xi_{i},\right.\\ \left.\gamma\in\widehat{\Gamma}^{+}(L_{0})\right\}$
and the set
$\left\{h\in\mathcal{O}_{K}/d\mathcal{O}_{K}\mid h\overline{h}\equiv-\ell\pmod{d}\textup{ and }\begin{pmatrix}(h\overline{h}+\ell)/d&h\\ \overline{h}&d\end{pmatrix}\sim f_{i}\right\}.$
Hence $n(\xi_{i};d)=r(f_{i};d)$. ∎
@@
Our next step is to relate the numbers $r(f_{i};d)$ to some representation numbers attached to the binary Hermitian form associated with the Hermitian matrix $f_{i}$.
@@
For each $i=1,\cdots,I_{\ell}$, we introduce the set
$P(f_{i},d):=\{(u,v)\in\mathcal{O}_{K}\times\mathcal{O}_{K}\,\,|\,\,f_{i}(u,v)=d,\,\,\,<u,v>=\mathcal{O}_{K}\},$
@@
where the notation $<u,v>$ is the ideal generated by $u$ and $v$ in $\mathcal{O}_{K}$. We moreover define the primitive representation numbers $p(f_{i},d):=\#P(f_{i},d)$. We define the set
$\mathcal{R}(d,-\ell):=\left\{h\in\mathcal{O}_{K}/d\mathcal{O}_{K}\mid h\overline{h}\equiv-\ell\pmod{d}\right\}$,
and following [EGM97, Definition 3.4], we define a map
$\varphi_{f_{i}}:P(f_{i},d)\rightarrow\mathcal{R}(d,-\ell),\,\,\,\varphi_{f_{i}}\left((u,v)\right):=h,$
@@
where for any $(u,v)\in P(f_{i},d)$, we pick $r,s\in\mathcal{O}_{K}$ such that $\begin{pmatrix}r&s\\ u&v\end{pmatrix}\in\textup{SL}_{2}(\mathcal{O}_{K})$ and set $h$ to be the entry
$\begin{pmatrix}r&s\\ u&v\end{pmatrix}f_{i}\begin{pmatrix}\bar{r}&\bar{u}\\ \bar{s}&\bar{v}\end{pmatrix}=\begin{pmatrix}*&h\\ \bar{h}&d\end{pmatrix}$.

### v2

@@ 10. Representation Numbers of Binary Hermitian Forms
The aim of this Section is to relate the expressions $n(\xi_{\mathfrak{m},i};d)$ of equation (6.3) with $\xi_{\mathfrak{m},i}\in L_{\mathfrak{m},0}^{+}[\ell,2^{-1}\mathbb{Z}]/\widehat{\Gamma}(L_{\mathfrak{m},0})$ to representation numbers of binary Hermitian forms.
We recall the map $f_{\omega}:(U,\psi)\longrightarrow(V_{0},\phi_{0})$ of Lemma 7.1. Since $h_{\mathfrak{m}}\in G_{\mathbb{Q}}^{*}$ and the map $f_{\omega}$ is an isometry, we have that there is an element $\beta_{\mathfrak{m}}\in\textup{SO}^{\psi}(U)$ such that $g_{\omega}(\beta_{\mathfrak{m}})=h_{\mathfrak{m}}$, where $g_{\omega}$ is the map (7.2). This means $f_{\omega}\beta_{\mathfrak{{m}}}f_{\omega}^{-1}=h_{\mathfrak{m}}$, or equivalently $f_{\omega}^{-1}h_{\mathfrak{m}}^{-1}v=\beta_{\mathfrak{m}}^{-1}f_{\omega}^{-1}v$ for all $v\in V_{0}$. Now, since the map $\tau$ is surjective, there is an element $\delta_{\mathfrak{m}}\in G^{+}(U)$ such that $\tau(\delta_{\mathfrak{m}})=\beta_{\mathfrak{m}}$. From the definition of the spinor norm, we have $\nu(\delta_{\mathfrak{m}})(\mathbb{Q}^{\times})^{2}=\sigma(\beta_{\mathfrak{m}})=\sigma(h_{\mathfrak{m}})=N(\mathfrak{m})(\mathbb{Q}^{\times})^{2}$, hence $\nu(\delta_{\mathfrak{m}})=N(\mathfrak{m})a^{2}$, with $a\in\mathbb{Q}^{\times}$. Let then $\delta_{\mathfrak{m}}^{\prime}:=a^{-1}\delta_{\mathfrak{m}}\in G^{+}(U)$. Then, we still have $\tau(\delta_{\mathfrak{m}}^{\prime})=\beta_{\mathfrak{m}}$
and $\nu(\delta_{\mathfrak{m}}^{\prime})=N(\mathfrak{m})$. If now $\alpha_{\mathfrak{m}}:=\Xi(\delta_{\mathfrak{m}}^{\prime})$, where $\Xi$ is the map of Proposition 7.4, with $\alpha_{\mathfrak{m}}\in\mathcal{G}$, then $\Xi((\delta_{\mathfrak{m}}^{\prime})^{-1})=\alpha_{\mathfrak{m}}^{-1}$, and we have
$\beta_{\mathfrak{m}}^{-1}u=\tau((\delta_{\mathfrak{m}}^{\prime})^{-1})u=\det(\alpha_{\mathfrak{m}})\alpha_{\mathfrak{m}}^{-1}u\overline{\alpha_{\mathfrak{m}}}^{-t},$
for all $u\in U$, using Proposition 7.4. We have the following Lemma.
Lemma 10.1.
The matrix $\alpha_{\mathfrak{m}}$ above satisfies $\det(\alpha_{\mathfrak{m}})=N(\mathfrak{m})$ and
$\alpha_{\mathfrak{m}}^{-1}\in\begin{pmatrix}\overline{\mathfrak{m}}^{-1}&\mathfrak{m}^{-1}\\ \overline{\mathfrak{m}}^{-1}&\mathfrak{m}^{-1}\end{pmatrix}$.
Proof.
The condition on the determinant follows from the fact that $\det(\alpha_{\mathfrak{m}})=\nu(\delta^{\prime}_{\mathfrak{m}})=N(\mathfrak{m})$ where $\delta^{\prime}_{\mathfrak{m}}\in G^{+}(U)$ is defined above. We now show the claim regarding the entries of the matrix.
We write $\alpha_{\mathfrak{m}}^{-1}=\begin{pmatrix}u&v\\ r&s\end{pmatrix}\in\textup{GL}_{2}(K)$. Since
$\Lambda_{\mathfrak{m},0}:=f^{-1}_{\omega}(L_{\mathfrak{m},0})=\begin{pmatrix}\mathbb{Z}&L_{\mathfrak{m}}\\ L_{\mathfrak{m}}&\mathbb{Z}\end{pmatrix}$,
and $L_{0}=h_{\mathfrak{m}}^{-1}L_{\mathfrak{m},0}$, by translating the actions as above, this means $\det(\alpha_{\mathfrak{m}})\alpha_{\mathfrak{m}}^{-1}\Lambda_{\mathfrak{m},0}\overline{\alpha_{m}}^{-t}=\Lambda$. Hence, we must have
$\begin{pmatrix}u&v\\ r&s\end{pmatrix}\begin{pmatrix}N(\mathfrak{m})&0\\ 0&0\end{pmatrix}\begin{pmatrix}\bar{u}&\bar{r}\\ \bar{v}&\bar{s}\end{pmatrix}=\begin{pmatrix}N(u)N(\mathfrak{m})&u\bar{r}N(\mathfrak{m})\\ \bar{u}rN(\mathfrak{m})&N(r)N(\mathfrak{m})\end{pmatrix}\in\begin{pmatrix}\mathbb{Z}&\mathcal{O}_{K}\\ \mathcal{O}_{K}&\mathbb{Z}\end{pmatrix}$.
That is $N(u)\in N(\mathfrak{m})^{-1}\mathbb{Z}$, $u\bar{r}\in N(\mathfrak{m})^{-1}\mathcal{O}_{K}$ and $N(r)\in N(\mathfrak{m})^{-1}\mathbb{Z}$.
A similar calculation with the matrix $\begin{pmatrix}0&0\\ 0&N(\mathfrak{m})\end{pmatrix}$ gives $N(v),N(s)\in N(\mathfrak{m})^{-1}\mathbb{Z}$ and $v\bar{s}\in N(\mathfrak{m})^{-1}\mathcal{O}_{K}$
Moreover we calculate (for an $h\in L_{\mathfrak{m}}=\frac{\bar{\mathfrak{m}}^{2}}{N(\mathfrak{m})}$),
$\begin{pmatrix}u&v\\ r&s\end{pmatrix}\begin{pmatrix}0&N(\mathfrak{m})h\\ N(\mathfrak{m})\bar{h}&0\end{pmatrix}\begin{pmatrix}\bar{u}&\bar{r}\\ \bar{v}&\bar{s}\end{pmatrix}=N(\mathfrak{m})\begin{pmatrix}v\bar{u}\bar{h}+u\bar{v}h&rv\bar{h}+u\bar{s}h\\ *&\bar{r}s\bar{h}+r\bar{s}h\end{pmatrix}\in\begin{pmatrix}\mathbb{Z}&\mathcal{O}_{K}\\ \mathcal{O}_{K}&\mathbb{Z}\end{pmatrix}$
That is, $N(\mathfrak{m})Tr(\bar{u}v\bar{h})\in\mathbb{Z}$, $N(\mathfrak{m})Tr(r\bar{s}h)\in\mathbb{Z}$ Equivalently, since $N(\mathfrak{m})h\in\bar{\mathfrak{m}}^{2}$ we have $\bar{u}v\in\mathfrak{d}^{-1}_{K}\mathfrak{m}^{-2}$ and $r\bar{s}\in\mathfrak{d}_{K}^{-1}\bar{\mathfrak{m}}^{-2}$. Since we are taking that $\mathfrak{d}_{K}$ and $\mathfrak{m}$ share no common factors, from the conditions $N(u),N(v)\in N(\mathfrak{m})^{-1}\mathbb{Z}$ and $\bar{u}v\in\mathfrak{d}^{-1}_{K}\mathfrak{m}^{-2}$ we have that $u\in\bar{\mathfrak{m}}^{-1}$ and $v\in\mathfrak{m}^{-1}$. Similarly we conclude that $r\in\bar{\mathfrak{m}}^{-1}$ and $s\in\mathfrak{m}^{-1}$. That is, the matrix
$\alpha^{-1}_{\mathfrak{m}}\in\begin{pmatrix}\bar{\mathfrak{m}}^{-1}&\mathfrak{m}^{-1}\\ \bar{\mathfrak{m}}^{-1}&\mathfrak{m}^{-1}\end{pmatrix}$$\qed$
Back to the expressions $n(\xi_{\mathfrak{m},i};d)$ of equation (6.3), we recall
@@
$n(\xi_{\mathfrak{m},i};d):=\#\left\{s\in\mathbb{Z}^{2}/dS_{\mathfrak{m}}\mathbb{Z}^{2}\mid D_{\mathfrak{m}}\equiv\frac{1}{2}q_{\mathfrak{m}}S_{\mathfrak{m}}^{-1}[s]\pmod{q_{\mathfrak{m}}d},\right.\\ \left.\left(\frac{\frac{1}{2}q_{\mathfrak{m}}S_{\mathfrak{m}}^{-1}[s]-D_{\mathfrak{m}}}{q_{\mathfrak{m}}d},A_{\mathfrak{m}}S_{\mathfrak{m}}^{-1}s,d\right)^{t}=\gamma\xi_{\mathfrak{m},i},\gamma\in\widehat{\Gamma}^{+}(L_{\mathfrak{m},0})\right\},$
where $D_{\mathfrak{m}}=-\ell q_{\mathfrak{m}}$, with $\ell$ a positive integer and $q_{\mathfrak{m}}$ the level of the lattice $L_{\mathfrak{m}}$.
Using Lemma 4.4, for any $\gamma\in\widehat{\Gamma}^{+}(L_{\mathfrak{m},0})$, we may write $\gamma=h_{\mathfrak{m}}\gamma^{\prime}h_{\mathfrak{m}}^{-1}$ with $\gamma^{\prime}\in\widehat{\Gamma}^{+}(L_{0})$, and so $\gamma\xi_{\mathfrak{m},i}=h_{\mathfrak{m}}\gamma^{\prime}\xi_{i}$, since $\xi_{\mathfrak{m},i}=h_{\mathfrak{m}}\xi_{i}$. Therefore, the above can be written as
(10.1) $n(\xi_{\mathfrak{m},i};d)=\#\left\{s\in\mathbb{Z}^{2}/dS_{\mathfrak{m}}\mathbb{Z}^{2}\mid D_{\mathfrak{m}}\equiv\frac{1}{2}q_{\mathfrak{m}}S_{\mathfrak{m}}^{-1}[s]\pmod{q_{\mathfrak{m}}d},\right.\\ \left.h_{\mathfrak{m}}^{-1}\left(\frac{\frac{1}{2}q_{\mathfrak{m}}S_{\mathfrak{m}}^{-1}[s]-D_{\mathfrak{m}}}{q_{\mathfrak{m}}d},A_{\mathfrak{m}}S_{\mathfrak{m}}^{-1}s,d\right)^{t}=\gamma\xi_{i},\ \gamma\in\widehat{\Gamma}^{+}(L_{0})\right\}.$
Putting together all the above, and with the matrix $\alpha_{\mathfrak{m}}\in\textup{GL}_{2}(K)$ with $\textup{det}(\alpha_{\mathfrak{m}})=N(\mathfrak{m})$ as above, we have that
(10.2) $f_{\omega}^{-1}\left(h_{\mathfrak{m}}^{-1}\left(\frac{\frac{1}{2}q_{\mathfrak{m}}S_{\mathfrak{m}}^{-1}[s]-D_{\mathfrak{m{}}}}{q_{\mathfrak{m}}d},A_{\mathfrak{m}}S_{\mathfrak{m}}^{-1}s,d\right)^{t}\right)=\det(\alpha_{\mathfrak{m}})\alpha^{-1}_{\mathfrak{m}}\begin{pmatrix}\frac{\frac{1}{2}q_{\mathfrak{m}}S_{\mathfrak{m}}^{-1}[s]-D_{\mathfrak{m{}}}}{q_{\mathfrak{m}}d}&h\\ \bar{h}&d&\end{pmatrix}\overline{\alpha_{\mathfrak{m}}}^{-t},$
where $h\in K$ corresponding to the vector $A_{\mathfrak{m}}S_{\mathfrak{m}}^{-1}s$, that is $h=\theta(A_{\mathfrak{m}}S_{\mathfrak{m}}^{-1}s)$, where $\theta:V\longrightarrow K$ is the isometry (4.2).
From Lemma 7.2 and Proposition 7.3, we have seen the identification
@@
induced by the map $f_{\omega}$. We set $f_{i}:=f_{\omega}^{-1}(\xi_{i})\in\Lambda^{+}[\ell,2^{-1}\mathbb{Z}]$, for $i=1,\ldots,h_{1}$. We have the following Lemma.
@@
Let $\Lambda=M_{2}(\mathcal{O}_{K})\cap U$. We have the following two statements:
@@
The first statement follows from [shimura2004arithmetic, Lemma 6.2, (3)], and the second statement is [shimura2006quadratic, Lemma 4.5]. ∎
By this Lemma, we have in particular that $f_{i}\in\Lambda$ since $\ell\in\mathbb{Z}$. Using (10.2) and the isomorphism $\Psi$ of Proposition 7.3, we then see that the equation
$h_{\mathfrak{m}}^{-1}\left(\frac{\frac{1}{2}q_{\mathfrak{m}}S_{\mathfrak{m}}^{-1}[s]-D_{\mathfrak{m}}}{q_{\mathfrak{m}}d},A_{\mathfrak{m}}S_{\mathfrak{m}}^{-1}s,d\right)^{t}=\gamma\xi_{i},\ \gamma\in\widehat{\Gamma}^{+}(L_{0})$
in (10.1) can be replaced with the condition that the matrix
$\det(\alpha_{\mathfrak{m}})\alpha^{-1}_{\mathfrak{m}}\begin{pmatrix}\frac{\frac{1}{2}q_{\mathfrak{m}}S_{\mathfrak{m}}^{-1}[s]-D_{\mathfrak{m}}}{q_{\mathfrak{m}}d}&h\\ \bar{h}&d&\end{pmatrix}\overline{\alpha_{\mathfrak{m}}}^{-t}\sim f_{i},$
where the notation $X\sim Y$ for two $2\times 2$ matrices $X$ and $Y$ means that there is a matrix $A\in\textup{SL}_{2}(\mathcal{O}_{K})$ such that $AX\overline{A}^{t}=Y$. Let now
(10.3) $\mathcal{N}(\xi_{\mathfrak{m},i};d):=\left\{s\in\mathbb{Z}^{2}/dS_{\mathfrak{m}}\mathbb{Z}^{2}\mid D\equiv\frac{1}{2}q_{\mathfrak{m}}S_{\mathfrak{m}}^{-1}[s]\pmod{q_{\mathfrak{m}}d},\,\,\alpha^{-1}_{\mathfrak{m}}A_{s}\overline{\alpha_{\mathfrak{m}}}^{-t}\sim f_{i}\right\},$
where
(10.4) $A_{s}=\begin{pmatrix}\frac{\frac{1}{2}q_{\mathfrak{m}}S_{\mathfrak{m}}^{-1}[s]-D}{q_{\mathfrak{m}}d}N(\mathfrak{m})&h_{s}N(\mathfrak{m})\\ \bar{h_{s}}N(\mathfrak{m})&dN(\mathfrak{m})&\end{pmatrix}$,
and $h_{s}:=\theta(A_{\mathfrak{m}}S_{\mathfrak{m}}^{-1}s)\in K$. We then have $n(\xi_{\mathfrak{m},i};d):=\#\mathcal{N}(\xi_{\mathfrak{m},i};d)$.
For a $d\in\mathbb{N}$, $f_{i}\in\Lambda$ with $\textup{det}(f_{i})=\ell$, we set
(10.5) $\mathcal{R}(f_{i},\mathfrak{m},d):=\left\{h\in\mathfrak{m}^{2}/d\mathfrak{m}^{2}\mid h\overline{h}\equiv-\ell N(\mathfrak{m})^{2}\pmod{dN(\mathfrak{m})^{2}}\right.\\ \left.\textup{ and }\exists\sigma\in\begin{pmatrix}\mathfrak{m}&\mathfrak{m}\\ \overline{\mathfrak{m}}&\overline{\mathfrak{m}}\end{pmatrix}\textup{ with }\det\sigma=N(\mathfrak{m})\textup{ such that }\right.\\ \left.\sigma f_{i}\sigma^{*}\sim\begin{pmatrix}\frac{(h\overline{h}+\ell N(\mathfrak{m})^{2})}{dN(\mathfrak{m})^{2}}N(\mathfrak{m})&h\\ \overline{h}&dN(\mathfrak{m})\end{pmatrix}\right\}$,
Let then $r(f_{i},\mathfrak{m},d):=\#\mathcal{R}(f_{i},\mathfrak{m},d)$. We now prove the following Proposition.
@@
Let $\ell\in\mathbb{Z}_{>0}$ and $\xi_{i}\in L_{0}^{+}[\ell,2^{-1}\mathbb{Z}]$. Recall $f_{i}:=f_{\omega}^{-1}(\xi_{i})\in\Lambda^{+}[\ell,2^{-1}\mathbb{Z}]$. Then, for any $d\in\mathbb{N}$, we have $n(\xi_{\mathfrak{m},i};d)=r(f_{i},\mathfrak{m},d)$.
@@
We consider the map $\mathcal{N}(\xi_{\mathfrak{m},i},d)\longrightarrow\mathcal{R}(f_{i},\overline{\mathfrak{m}},d)$ given by $s\longmapsto h_{s}:=\theta(A_{\mathfrak{m}}S_{\mathfrak{m}}^{-1}s)N(\mathfrak{m})$.
Well-defined: First of all, if $s\in\mathcal{N}(\xi_{\mathfrak{m},i},d)$, then we consider the vector
$v:=\left(\frac{\frac{1}{2}q_{\mathfrak{m}}S_{\mathfrak{m}}^{-1}[s]-D}{q_{\mathfrak{m}}d},A_{\mathfrak{m}}S_{\mathfrak{m}}^{-1}s,d\right).$
We observe that $\phi_{0}[v]=$-$\ell\in\mathbb{Z}$ and because of the conditions, we see that $v\in L_{\mathfrak{m},0}^{*}$, and $L_{\mathfrak{m},0}$ is $\mathbb{Z}$-maximal. This, by [shimura2004arithmetic, Lemma 6.2, (3)], in turn implies $v\in L_{\mathfrak{m},0}$, which shows $A_{\mathfrak{m}}S_{\mathfrak{m}}^{-1}s\in L_{\mathfrak{m}}$.
This now shows that $\theta(A_{\mathfrak{m}}S_{\mathfrak{m}}^{-1}s)\in\overline{\mathfrak{m}}^{2}/N(\mathfrak{m})$, and hence $h_{s}\in\overline{\mathfrak{m}}^{2}$. Moreover, since $\theta$ is an isometry,
$h_{s}\overline{h}_{s}=N_{K/\mathbb{Q}}(h_{s})=N_{K/\mathbb{Q}}(\theta(A_{\mathfrak{m}}S_{\mathfrak{m}}^{-1}s)N(\mathfrak{m}))=N(\mathfrak{m})^{2}\frac{1}{2}S[A_{\mathfrak{m}}S_{\mathfrak{m}}^{-1}s]=N(\mathfrak{m})^{2}\frac{1}{2}S_{\mathfrak{m}}^{-1}[s].$
Hence $h_{s}\overline{h}_{s}+\ell N(\mathfrak{m})^{2}=N(\mathfrak{m})^{2}\left(\frac{1}{2}S_{\mathfrak{m}}^{-1}[s]+\ell\right)\equiv 0\pmod{dN(\mathfrak{m})^{2}}$, because $D\equiv\frac{1}{2}q_{\mathfrak{m}}S_{\mathfrak{m}}^{-1}[s]\pmod{q_{\mathfrak{m}}d}$ implies $\frac{1}{2}S_{\mathfrak{m}}^{-1}[s]+\ell\equiv 0\pmod{d}$, after replacing $D=-q_{\mathfrak{m}}\ell$.
Assume now $s\equiv t\pmod{dS_{\mathfrak{m}}\mathbb{Z}^{2}}$ and write $s-t=dS_{\mathfrak{m}}u$, with $u\in\mathbb{Z}^{2}$. We then have
$h_{s}-h_{t}=N(\mathfrak{m})\theta(A_{\mathfrak{m}}S_{\mathfrak{m}}^{-1}(s-t))=dN(\mathfrak{m})\theta(A_{\mathfrak{m}}u),$
and since $L_{\mathfrak{m}}=A_{\mathfrak{m}}L$, we have $\theta(A_{\mathfrak{m}}u)\in\overline{\mathfrak{m}}^{2}/N(\mathfrak{m})$. This shows $h_{s}-h_{t}\in d\overline{\mathfrak{m}}^{2}$.
Finally, $\begin{pmatrix}\frac{(h_{s}\overline{h_{s}}+\ell N(\mathfrak{m})^{2})}{dN(\mathfrak{m})^{2}}N(\mathfrak{m})&h_{s}\\ \overline{h_{s}}&dN(\mathfrak{m})\end{pmatrix}=A_{s}$ (see (10.4)), and the last property follows from the property of $A_{s}$ and the properties of $\alpha_{\mathfrak{m}}$.
Injective: Assume now $h_{s}=h_{t}$ in $\mathcal{R}(f_{i},\overline{\mathfrak{m}},d)$. This means that there is some $a\in\overline{\mathfrak{m}}^{2}$ such that $h_{s}-h_{t}=da$. By the calculations above, we have
$h_{s}-h_{t}=(N\mathfrak{m})\theta(A_{\mathfrak{m}}S_{\mathfrak{m}}^{-1}(s-t)),$
which shows
$s-t=N(\mathfrak{m})^{-1}S_{\mathfrak{m}}A_{\mathfrak{m}}^{-1}\theta^{-1}(h_{s}-h_{t})=dS_{\mathfrak{m}}A_{\mathfrak{m}}^{-1}\theta^{-1}(N(\mathfrak{m})^{-1}a)\in dS_{\mathfrak{m}}\mathbb{Z}^{2},$
because $\theta(N(\mathfrak{m})^{-1}a)\in L_{\mathfrak{m}}$, since $N(\mathfrak{m})^{-1}a\in\overline{\mathfrak{m}}^{2}/N(\mathfrak{m})$ and therefore $A_{\mathfrak{m}}^{-1}L_{\mathfrak{m}}\in L=\mathbb{Z}^{2}$.
Surjective: If now $h\in\mathcal{R}(f_{i},\overline{\mathfrak{m}},d)$, we can choose $s:=S_{\mathfrak{m}}A_{\mathfrak{m}}^{-1}\theta^{-1}(N(\mathfrak{m})^{-1}h)\in\mathbb{Z}^{2}$, since $N(\mathfrak{m})^{-1}h\in\overline{\mathfrak{m}}^{2}/N(\mathfrak{m})$ and hence $\theta^{-1}(N(\mathfrak{m})^{-1}h)\in L_{\mathfrak{m}}$, hence $s\in S_{\mathfrak{m}}\mathbb{Z}^{2}\subset\mathbb{Z}^{2}$. ∎
@@
Our next step is to relate the numbers $r(f_{i},\overline{\mathfrak{m}},d)$ to some representation numbers attached to the binary Hermitian form associated with the Hermitian matrix $f_{i}$.
@@
For each $i=1,\cdots,I_{\ell}$, we introduce the set of $\mathfrak{m}$-primitive representations of $d$ by $f_{i}$ (see [Elstrodt, Definition 3.1])
$P(f_{i},\mathfrak{m},d):=\{(u,v)\in\mathfrak{m}\times\mathfrak{m}\,\,|\,\,f_{i}(u,v)=N(\mathfrak{m})d,\,\,\,<u,v>=\mathfrak{m}\},$
@@
where the notation $<u,v>$ is the fractional ideal generated by $u$ and $v$ in $K$. Moreover, we define the $\mathfrak{m}$-primitive representation numbers $p(f_{i},\mathfrak{m},d):=\#P(f_{i},\mathfrak{m},d)$. Finally, in addition to the set 10.5, we define
$\mathcal{R}(\mathfrak{m},d,-\ell):=\left\{h\in\mathfrak{m}^{2}/d\mathfrak{m}^{2}\mid h\overline{h}\equiv-\ell N(\mathfrak{m})^{2}\pmod{dN(\mathfrak{m})^{2}}\right)$.
Following [Elstrodt, Definition 3.4], we define a map
(10.6) $\varphi_{f_{i}}:P(f_{i},\mathfrak{m},d)\rightarrow\mathcal{R}(\mathfrak{m},d,-\ell),\,\,\,\varphi_{f_{i}}\left((u,v)\right):=h,$
@@
where for any $(u,v)\in P(f_{i},\mathfrak{m},d)$, we pick $r,s\in\overline{\mathfrak{m}}$ such that $\det\begin{pmatrix}r&s\\ u&v\end{pmatrix}=N(\mathfrak{m})$ and set $h$ to be the entry such that
$\begin{pmatrix}r&s\\ u&v\end{pmatrix}f_{i}\begin{pmatrix}\bar{r}&\bar{u}\\ \bar{s}&\bar{v}\end{pmatrix}=\begin{pmatrix}*&h\\ \bar{h}&dN(\mathfrak{m})\end{pmatrix}$.